### v1

# Flowing from Words to Pixels: A Framework for Cross-Modality Evolution

Abstract
Diffusion models, and their generalization, flow matching, have had a remarkable impact on the field of media generation. Here, the conventional approach is to learn the complex mapping from a simple source distribution of Gaussian noise to the target media distribution. For cross-modal tasks such as text-to-image generation, this same mapping from noise to image is learnt whilst including a conditioning mechanism in the model. One key and thus far relatively unexplored feature of flow matching is that, unlike Diffusion models, they are not constrained for the source distribution to be noise. Hence, in this paper, we propose a paradigm shift, and ask the question of whether we can instead train flow matching models to learn a direct mapping from the distribution of one modality to the distribution of another, thus obviating the need for both the noise distribution and conditioning mechanism. We present a general and simple framework, CrossFlow, for cross-modal flow matching. We show the importance of applying Variational Encoders to the input data, and introduce a method to enable Classifier-free guidance. Surprisingly, for text-to-image, CrossFlow with a vanilla transformer without cross attention slightly outperforms standard flow matching, and we show that it scales better with training steps and model size, while also allowing for interesting latent arithmetic which results in semantically meaningful edits in the output space. To demonstrate the generalizability of our approach, we also show that CrossFlow is on par with or outperforms the state-of-the-art for various cross-modal / intra-modal mapping tasks, viz. image captioning, depth estimation, and image super-resolution. We hope this paper contributes to accelerating progress in cross-modal media generation.
Figure 1: We propose CrossFlow, a general and simple framework that directly evolves one modality to another using flow matching with no additional conditioning. This is enabled using a vanilla transformer without cross-attention, achieving comparable performance with state-of-the-art models on (a) text-to-image generation, and (b) various other tasks, without requiring task specific architectures.
1 Introduction
Diffusion models have achieved remarkable success in generating images [17, 69, 59, 74, 73], videos [34, 8, 79, 7], audio [41, 53], and 3D content [65, 47], revolutionizing the field of generative AI. Recently, flow matching [50, 2, 55] has been proposed as a generalization of diffusion models, where models are trained to find an optimal transport probability path between a source noise distribution and the target data distribution. This approach offers simpler, straight-line trajectories compared to the complex, curved trajectories in diffusion paths. As a result, it has been rapidly adopted in the latest state-of-the-art image and video generation models, including LDMs [22] and Movie Gen [64].
Both diffusion and flow-based models are typically trained to learn the mapping from noise to the target distribution. For cross-modal generation tasks such as text-to-image [73, 10], this same mapping from noise to the target modality distribution (i.e. the images) is learnt whilst adding a conditioning mechanism for the conditioning modality (i.e. the text) such as cross-attention. Unlike denoising diffusion models [33, 82], one relatively unexplored feature of flow matching models is that they are not constrained for the source distribution to be Gaussian noise; instead, the source distribution could be one that is correlated with the target distribution. Compared to noise, learning a mapping from such a distribution should intuitively be easier for the model because it has to learn shorter and more efficient probability paths. A question remains however as to what this correlated source distribution could be.
Interestingly, due to the information redundancy between different modalities arising from the same data point, for cross-modal generation tasks, the provided conditioning (e.g. the text in text-to-image) resembles such data that is correlated with the target distribution (e.g. the images). Hence, in this paper, we propose a paradigm shift for cross-modal generation, and ask the question of whether we can instead train flow matching models to learn a direct mapping from the distribution of one modality to the distribution of another, hence obviating the need for both the noise distribution and any conditioning mechanism.
Despite the exciting theoretical motivation, there are several key challenges in practice. First, both diffusion and flow-based models require the source and target distributions to be of the same shape; a requirement that is not satisfied for data distributions from different modalities. Secondly, state-of-the-art methods heavily rely on Classifier-free guidance (CFG) [32] for improved generation quality; a method that is not compatible with cross-modal flow matching due to the lack of a conditioning mechanism to turn on/off since the conditioning information instead lies within the source data. As a result, prior work [29, 55, 1] targets the simple setting of mapping between two similar intra-modal distributions, such as human faces to cat faces [55].
In this work, we present key architecture design solutions for overcoming these challenges: First, we employ a Variational Encoder for encoding the source modality data distribution to the same shape as the target modality, and show that the resulting regularization in the source distribution is essential for generation performance. Secondly, we enable CFG in cross-modal flow matching through the introduction of a binary conditioning indicator during training, and demonstrate the quantative benefits of this approach compared to alternative CFG methods. We present CrossFlow; a general framework for mapping between two different modalities without the need for any conditioning mechanism or noise distribution. Typically, different cross-modal generation tasks require task-specific architectural and training modifications, but CrossFlow works for different tasks without any such changes.
Using the ubiquitous albeit challenging text-to-image (T2I) generation task as our primary setting, we show the significant result that CrossFlow outperforms commonly used flow matching baselines, given the same training data, model size, and training budget, all without requiring any cross-attention layers. CrossFlow exhibits improved scaling behavior over standard flow matching using cross-attention when scaling training steps or model size, and is also compatible with a variety of Large Language Models (LLMs), including CLIP [66], T5 [67], and Llama3 [19]. Additionally we demonstrate that since our approach encodes the source distribution into a regularized continuous space with semantic structure, CrossFlow enables exciting new latent arithmetic for the text-to-image task, e.g., $\mathcal{L}$(“A dog with a hat”) + $\mathcal{L}$(“Sunglasses”) – $\mathcal{L}$(“A hat”) creates an image of a dog wearing sunglasses without a hat.
We demonstrate the general-purpose nature of CrossFlow on various cross-modal/intra-modal tasks: image-to-text (image captioning), image-to-depth (depth estimation), and low-resolution to high-resolution image (super-resolution). we show that our CrossFlow approach achieves comparable or even superior performance to various state-of-the-art methods on all three tasks. Results are shown in Fig. 1. We hope that this paper contributes to accelerating the progress in cross-modal media generation.
2 Related Work
Diffusion models and Rectified Flow. Starting from Gaussian noise, diffusion [80, 33] and score-based [36, 81] generative models progressively approximate the reverse ODE of a stochastic forward process to generate data. These models have driven significant advances across various domains, particularly in high-fidelity image [17, 35, 62, 5, 54], video [34, 8, 79, 7, 64], and 3D generation [65, 47]. The rapid development of generative models in recent years demonstrates the effectiveness and stability of diffusion models in modeling complex real-world distributions. Recently, rectified flow models [50, 2, 55], such as flow matching, have been proposed to improve the generative process by enabling a transport map between two distributions via ODE. They enable faster training and sampling by avoiding complex probability flow ODEs.
Directly bridging distributions. Flow Matching theoretically allows for arbitrary distributions as the source distribution, which can then be used for direct evolution. Various approaches have been proposed in this direction, such as InterFlow [1], $\alpha$-blending [29], and Schrödinger Bridge [52, 77, 95, 51, 15, 83, 84]. These methods provide important theoretical support for using ODE-based methods to bridge two arbitrary distributions. However, they are still limited to the similar distributions from the same domain, such as image-to-image translation (e.g., faces-to-faces [55, 95] or sketches-to-images [52]). As a step forward, CrossFlow focuses on learning the mapping between data distributions arising from even different modalities.
Text-to-Image generation. Text-to-image generation [68, 59, 74, 73, 69, 10, 22, 14] has witnessed rapid advancements with the advent of diffusion and subsequently flow matching models. This task bridges two complex and important domains: language and vision. To tackle this challenge, existing methods often incorporate text encoders such as LLMs into the diffusion model through additional conditioning mechanisms, with cross-attention being the most widely adopted [22, 64]. However, these approaches increase modeling complexity and require additional parameters. We demonstrate that CrossFlow improves over standard flow matching with better scaling characteristics, and is comparable to prior work, despite a simpler architecture.
Cross-modal / intra-modal mapping. Various tasks can be framed as cross-modal/intra-modal mapping problems, including image captioning [44, 93, 24, 43, 28, 96, 58], depth estimation [90, 70, 39, 6, 45, 46, 18], and image super-resolution [75, 23]. However, due to the significant differences between modalities or distributions, previous methods have typically relied on task-specific designs. For example, Bit Diffusion [12] encodes text into binary bits and uses a diffusion model with self-conditioning for captioning. Flow-based super-resolution models, such as CFM [23], still require the low-resolution image as extra conditioning, and also add Gaussian noise to the input. In contrast, our CrossFlow uses the same unified framework across all these tasks without extra conditioning or noise.
3 Preliminaries
Figure 2: CrossFlow Architecture. CrossFlow enables direct evolution between two different modalities. Taking text-to-image generation as an example, our T2I model comprises two main components: a Text Variational Encoder and a standard flow matching model. At inference time, we utilize the Text Variational Encoder to extract the text latent $z_{0}\in\mathbb{R}^{h\times w\times c}$ from text embedding $x\in\mathbb{R}^{n\times d}$ produced by any language model. Then we directly evolve this text latent into the image space to generate image latent $z_{1}\in\mathbb{R}^{h\times w\times c}$.
Flow Matching. We consider a generative model that defines a mapping between samples $z_{0}$ from a source distribution $p_{0}$ to samples $z_{1}$ of a target distribution $p_{1}$ via the ordinary differential equation (ODE): $dz_{t}=v_{\theta}(z_{t},t)dt$. Here, $v_{\theta}$ is the velocity parameterized by the weights $\theta$ of a neural network, and $t\in[0,1]$ is the time-step.
Flow Matching [50, 2, 55] defines the forward process using the optimal transport path, i.e.,
$z_{t}=tz_{1}+(1-(1-\sigma_{min})t)z_{0}$ (1)
where $\sigma_{min}=10^{-5}$. The ground truth velocity can be derived as:
$\hat{v}_{t}=\frac{dz_{t}}{dt}=z_{1}-(1-\sigma_{min})z_{0}$ (2)
To achieve this, a network $v_{\theta}(z_{t},t)$ is trained to predict velocity by minimizing the mean squared error (MSE) between its output and the target $\hat{v}_{t}$. This constructs a continuous path between $z_{0}$ and $z_{1}$ at any time-step $t\in[0,1]$.
As discussed earlier, flow matching enables evolving a sample $z_{1}$ from an arbitrary source distribution $p_{0}$. But prior work [22, 64] has typically relied on starting from a simple Gaussian noise sample $z_{0}\sim\mathcal{N}(0,1)$, and computing the velocity with additional condition $c$ incorporated through various methods, e.g., cross-attention [22, 64], channel-wise concatenation [27].
Classifier-free guidance. CFG [32] is a broadly used technique that enhances sample quality in conditional generative models by jointly training a single model on conditional and unconditional objectives. This is achieved through randomly dropping the condition $c$ during training with a certain probability $p$. Sampling is performed by extrapolating between conditional and unconditional denoising $v_{\theta}(z_{t},c)$ and $v_{\theta}(z_{t})$ with a scaling factor $\omega$:
$\tilde{v}_{\theta}(z_{t},c)=\omega v_{\theta}(z_{t},c)+(1-\omega)v_{\theta}(z_{t})$ (3)
It significantly improves the generation quality and fidelity by guiding the samples towards higher likelihood of the condition $c$, which plays a crucial role in state-of-the-art media generation models [22, 10, 69, 64].
4 CrossFlow
In this section, we discuss the various components of our approach: a Variational Encoder (VE) to encode the inputs in Sec. 4.1, using flow matching to evolve from the source to the target distribution in Sec. 4.2, and finally, applying CFG in this setting for improving quality and fidelity in Sec. 4.3.
4.1 Variational Encoder for Encoding Inputs
Flow matching models require the source distribution $p_{0}$ to have the same shape as the target distribution $p_{1}$. In particular, given an input $x$, we need to convert it to the source latent $z_{0}$, which has the same shape as the target latent $z_{1}$. An intuitive solution is to use an encoder $\mathcal{E}$ to convert $x$ to $z_{0}$, i.e., $z_{0}=\mathcal{E}(x)$, which can preserve most of the input information as shown in Tab. 9. However, directly evolving from $\mathcal{E}(x)$ to $z_{1}$ is problematic. We find that it is essential to formulate $z_{0}$ as a regularized distribution for the source in order for flow matching to work well. To address this, we propose using a VE to convert $x$ to $z_{0}$. Formally, instead of directly predicting $z_{0}$, we predict its mean $\bar{\mu}_{z_{0}}$ and variance $\bar{\sigma}_{z_{0}}$, and then sample the latent $z_{0}\sim\mathcal{N}(\bar{\mu}_{z_{0}},\bar{\sigma}_{z_{0}}^{2})$. This enables us to convert the given input $x$ into latent $z_{0}$ with a regularized distribution, which can then be gradually evolved into the target distribution $z_{1}$ with flow matching.
The VE can be trained with a standard Variational Autoencoding objective (VAE) [40] comprising of an encoding loss and the KL-divergence loss. For the encoding loss, the VE is trained to minimize a loss between the output $z_{0}$ and a target $\hat{z}$. For a VAE this loss would be a reconstruction loss like MSE between the input $x$ and the decoder $\mathcal{D}$’s output, $\texttt{MSE}(\mathcal{D}(z_{0}),x)$. But since we simply need a encoder and not an autoencoder, we don’t restrict ourselves to a VAE.
4.2 Training CrossFlow
For each training sample, we start with an input-target pair $(x,z_{1})$. We apply the VE to $x$ to encode it to a latent $z_{0}$ with the same shape as $z_{1}$. Next, we employ a transformer model $v_{\theta}$ trained for flow matching as per Equations 1 and 2. The VE can be trained prior to training $v_{\theta}$ or concurrently. We show in Sec. 5.2 that jointly training the Variational Encoder with flow matching results in improved performance.
Specifically, we jointly train the VE with the flow matching model using a sum of flow matching MSE loss $L_{FM}$, and the losses for Variational Encoder training (encoding loss $L_{Enc}$ and KL-divergence loss $L_{KL}$):
$L$ $=L_{FM}+L_{Enc}+\lambda L_{KL}$
$=\texttt{MSE}(v_{\theta}(z_{t},t),\hat{v})+\texttt{Enc}(z_{0},\hat{z})$
$\qquad+\lambda\texttt{KL}(\mathcal{N}(\bar{\mu}_{z_{0}},\bar{\sigma}_{z_{0}}^{2})||\mathcal{N}(0,1))$ (4)
where $\lambda$ is the weight of KL-divergence loss. Please find more details in Sec. A.1.
4.3 Classifier-Free Guidance with an Indicator
CFG [32] has become the standard low-temperature sampling method for enhancing multi-modal alignment and improving quality. However, it can only be applied to generation methods that accept an additional conditioning input $c$, since the guidance signal relies on the difference between conditional and unconditional predictions $v_{\theta}(z_{t},c)$ and $v_{\theta}(z_{t})$. Recently, Autoguidance (AG) [38] has been introduced as a method to enhance both conditional and unconditional generation, by leveraging a smaller and less-trained “bad model” as guidance. Nevertheless, it falls short of standard CFG in terms of performance. Moreover, AG requires training a separate bad model, and its performance varies dramatically based on the choice of the bad model. The search space can be restricted by using an under-trained version of the same model, but this affects performance, and also is prohibitive for large models since it requires loading two large models in memory for inference.
We instead aim to support CFG for CrossFlow, which is as accessible and performant as CFG is for standard flow matching. To enable CFG without the presence of an explicit conditioning input $c$, we introduce CFG with an indicator. Specifically, our model is of the form $v_{\theta}(z_{t},1_{c})$, where $1_{c}\in\{0,1\}$ is an indicator to specify conditional vs. unconditional generation. The model evolves from $z_{0}$ to $z_{1}$ when $11_{c}=$ 1, and from $z_{0}$ to $z_{1}^{uc}$ when $01_{c}=$ 0, where $z_{1}^{uc}$ represents any sample from the target distribution $p_{1}$ other than $z_{1}$. During training, we employ two learnable parameters, $g^{c}$ and $g^{uc}$, corresponding to conditional and unconditional generation, respectively. Depending on $1_{c}$, the appropriate learnable parameter is concatenated with the transformer input tokens along sequence dimension. We randomly sample the indicator with an unconditional rate of $10\%$, as per standard practice. The insight behind the CFG indicator is similar to that of standard CFG. In this approach, $v_{\theta}(z_{t},1)$ is trained to map $z_{0}$ to a specific region of the target manifold, while $v_{\theta}(z_{t},0)$ is trained to map $z_{0}$ to the entire target manifold.
4.4 Flowing from Text to Image
Now, we consider text-to-image generation as the archetypal task to leverage CrossFlow. We start with the input text embedding $x\in\mathbb{R}^{n\times d}$ with token length $n$ and dimension $d$, and use our Text VE to extract the corresponding text latent $z_{0}\sim\mathcal{N}(\bar{\mu}_{x},\bar{\sigma}_{x}^{2})$. While our approach is agnostic to pixel vs. latent image generation, we consider image generation in the latent space for efficiency, and leverage a pre-trained VAE to obtain the image latent from the input image I, which serves as our target $z_{1}$. Then, we employ the vanilla flow matching [50] model to predict $v(z_{t},t)$ between $z_{0}$ and $z_{1}$. The pipeline for performing text-to-image generation with CrossFlow is illustrated in Fig. 2. We discuss how to train a performant Text Variational Encoder next.
4.4.1 Text Variational Encoder
Training the Text VE is challenging, as this involves compressing the text embeddings to small latent space (e.g., $77\times 768$ CLIP tokens to $4\times 32\times 32$ image latents for $256$px generation, 14.4×14.4$\times$14.4 × compression). We explore various methods to train VEs for CrossFlow. The straightforward approach is to simply train a VAE with a MSE reconstruction loss. While this approach achieves very low reconstruction errors, we find that it does not capture semantic concepts well, leading to sub-optimal image generations.
Contrastive loss. We explore contrastive losses, which produce representations with strong semantic understanding when training on samples within the same modality [60, 11] and on different modality pairs [66]. To produce the contrastive targets for the VE, we either use the input text embedding $x$ (text-text contrastive), or the paired image $I$ for the text (image-text contrastive). Given the target, we employ a simple encoder to project it into a feature space with the same shape as $z_{0}$, resulting in a representation denoted as $\hat{z}$. We then encourage semantic similarity between $z_{0}$ and $\hat{z}$ using the contrastive CLIP loss [66]. During training, the batch-wise contrastive loss is computed as $L_{Enc}=\texttt{CLIP}(z_{0},\hat{z})$. We ablate this choice in Sec. 5.2 and find that contrastive loss works significantly better than the VAE reconstruction loss, with the image-text loss working slightly better than the text-text loss.
5 Experiments
We first evaluate CrossFlow on text-to-image generation, demonstrate its scalability, and showcase some interesting applications with latent arithmetic in Sec. 5.1. Then, we ablate our main design decisions through ablation studies in Sec. 5.2. Finally, we further explore CrossFlow’s performance on three distinct tasks: image captioning, monocular depth estimation, and image super-resolution in Sec. 5.3.
5.1 Text-to-Image Generation
Experimental setup. Comparing T2I models scientifically is challenging due to the diverse datasets used for training, which often include proprietary data, and varying training conditions. In addition, our proposed method represents a new paradigm for utilizing diffusion models, distinct from the previous T2I approaches. Therefore, we primarily compare our model with the widely used “standard flow matching baseline” that starts from noise and leverages text cross-attention. For a fair comparison, we use the exact same codebase, training recipe, dataset, and budget to train both CrossFlow and the baseline. Note that the baseline requires a cross-attention layer after each self-attention layer, whereas our model only relies on self-attention layers, resulting in fewer parameters for the same number of layers. To account for this, we adjust the number of layers to ensure that both models have similar model sizes. For both methods, we use a grid search to find the optimal CFG scale. We also compare CrossFlow with state-of-the-art T2I models to demonstrate that our approach is competitive with those established methods.
Architecture. Our model enables the use of vanilla Transformer [86] with self-attention layers and feed-forward layers. We adopt DiMR [54] as the base architecture for the flow matching models, a variant of Diffusion Transformer (DiT) [62] which replaces the parameter-heavy MLP in adaLN-Zero with a more parameter-efficient Time-Dependent Layer Normalization. For the Text VE, we employ stacked Transformer blocks, followed by a linear layer to project the output into the target shape.
Method #Params (B) #Steps (K) FID $\downarrow$ CLIP $\uparrow$
Standard FM (Baseline) 1.04 300 10.79 0.29
CrossFlow (Ours) 0.95 300 10.13 0.29
Table 1: Comparison between our CrossFlow and standard flow matching with cross-attention. Both models are trained with the same settings. We find that our model slightly outperforms standard flow matching baseline in terms of zero-shot FID-30K and achieves comparable performance on the CLIP score.
Training details. We use a proprietary dataset with about $350$M image-text pairs to train both CrossFlow and our ablations. Our text encoder is based on CLIP [66] with a fixed sequence length of $77$ text tokens. We use a pre-trained and frozen VAE from LDM [73] to extract image latents. All T2I models are trained using the same settings: an image resolution of $256\times 256$, a batch size of $1024$, a base learning rate of $1\times 10^{-4}$ with $5000$ warm-up steps, and an AdamW optimizer [57] with $\beta_{1}=\beta_{2}=0.9$ and a weight decay of $0.03$. We train our largest model ($0.95$B) on $256\times 256$ for $600$K iterations, then finetune it on $512\times 512$ for an additional $240$K iterations to generate higher-resolution images.
701503005001000666610101010141414141818181822222222Model parameters (millions)FID-30K
50100150200250300121212121313131314141414151515151616161617171717Training steps (thousands)FID-30KStandard FMCrossFlow
Figure 3: Performance vs. Model Parameters and Iterations. We compare the baseline of starting from noise with text cross-attention with CrossFlow, while controlling for data, model size and training steps. Left: Larger models are able to exploit the cross-modality connection better. Right: CrossFlow needs more steps to converge, but converges to better final performance. Overall, CrossFlow scales better than the baseline and can serve as the framework for future media generation models.
Evaluation metrics. We evaluate all models on the COCO validation set [49] and report FID [31] and CLIP score [66, 30]. Following previous works, we report zero-shot FID-30K, where 30K prompts are randomly sampled from the validation set, and the generated images are compared to reference images from the full validation set. Additionally, we also evaluate our models on GenEval benchmark as it exhibits strong alignment with human judgment [26].
5.1.1 CrossFlow vs. Standard Flow Matching
We compare our CrossFlow with widely used cross-attention baseline in Tab. 1. Both models are trained and tested under the same settings. The results show that CrossFlow achieves comparable performance, with slightly better zero-shot FID-30K compared with widely used flow matching baselines with cross-attention.
Scaling characteristics. We investigate the scalability of CrossFlow in Fig. 3 and compare it with standard flow matching. We train both approaches across 5 different model sizes, ranging from $70$M to $1$B parameters, with the same training settings, for $300$K iterations. At smaller scales, CrossFlow underperforms the baseline, likely due to the lack of sufficient parameters to model the complex relationships between two modalities. But excitingly, as the model size increases, the zero-shot FID-30K improves more for our approach. Next, we evaluate the effect of varying the training iterations. We notice similarly that CrossFlow improves more as we increase training iterations.
While CrossFlow initially underperforms standard flow matching at small scales, increasing the model size and training iterations improves it significantly, even enabling it to surpass standard flow matching. We attribute this to the fact that CrossFlow generates images by directly evolving from the source distribution where different sub-regions correspond to different semantics. In contrast, standard flow matching may generate the same semantics from the entire source distribution, while exploiting the inductive biases afforded by text cross-attention. Ultimately, this works in favor of CrossFlow, as the learnt cross-modal paths and fewer inductive biases result in improved scaling characteristics with both model size and training iterations.
Method #Params. FID-30K $\downarrow$ GenEval $\uparrow$
zero-shot score
DALL·E [68] 12.0B 27.50 -
GLIDE [59] 5.0B 12.24 -
LDM [73] 1.4B 12.63 -
DALL·E 2 [69] 6.5B 10.39 0.52
LDMv1.5 [73] 0.9B 9.62 0.43
Imagen [74] 3.0B 7.27 -
RAPHAEL [88] 3.0B 6.61 -
PixArt-$\alpha$ [10] 0.6B 7.32 0.48
LDMv3 ($512^{2}$[22] 8.0B - 0.68
CrossFlow 0.95B 9.63 0.55
Table 2: Comparison with recent T2I models. For GenEval, we report the overall score here and provide task-specific scores in Sec. B.1. CrossFlow achieves comparable performance with state-of-the-art T2I models by directly evolving text into images.
5.1.2 State-of-the-art Comparison
Finally, we compare CrossFlow with state-of-the-art text-to-image models and report results in Tab. 2. We achieve a zero-shot FID-30K of $9.63$ on COCO, and a GenEval score of $0.55$, demonstrating performance comparable with the state-of-the-art. Note that our model uses only 630 A100 GPU days for training, whereas other methods like DALL·E 2 [69] typically require thousands of A100 GPU days. These results suggest that CrossFlow is a simple and promising direction for state-of-the-art media generation.
Figure 4: CrossFlow provides visually smooth interpolations in the latent space. We show images generated by linear interpolation between the first (left) and second (right) text latents. CrossFlow enables visually smooth transformations of object direction, composite colors, shapes, background scenes, and even object categories. Please zoom in for better visualization. For brevity, we display only 7 interpolating images here; additional interpolating images can be found in Appendix C (Fig. 10 and Fig. 11).
Figure 5: CrossFlow allows arithmetic in text latent space. Using the Text Variational Encoder (VE), we first map the input text into the latent space $z_{0}$. Arithmetic operations are then performed in this latent space, and the resulting latent representation is used to generate the corresponding image. The latent code $z_{0}$ used to generate each image is provided at the bottom.
5.1.3 Arithmetic Operations in Latent Space
Unlike previous diffusion or flow matching models, CrossFlow offers a unique property: arithmetic operations in the input latent space translate to similar operations in the output space. This is made possible since CrossFlow transforms the source space (i.e., the text latent space for T2I) into a regularized continuous space, where a uniform representation shape is shared across all texts. We showcase two examples of this, latent interpolation, and latent arithmetic. For linear interpolation, we use the Text Variational Encoder to generate text latents from two different text inputs, and then interpolate between them to produce images. As shown in Fig. 4, CrossFlow enables visually smooth linear interpolations, even between disparate prompts. Next, we showcase arithmetic operations in Fig. 5, in which we apply addition and subtraction in the text latent space, and find that the resulting images exhibit corresponding semantic modifications to the original image. This demonstrates that CrossFlow formulates meaningful and well-structured semantic paths between the source and the target distributions, providing interesting capabilities and more control compared to standard flow matching approaches.
5.2 Ablation Study
We conduct various ablation experiments to verify the effectiveness of the proposed designs in Tab. 3.
Text encoder FID $\downarrow$ CLIP $\uparrow$
Encoder 66.65 0.20
Encoder + noise 59.91 0.21
Variational Encoder 40.78 0.23
Loss FID $\downarrow$ CLIP $\uparrow$
T-T Recon. 40.78 0.23
T-T Contrast. 34.67 0.24
I-T Contrast. 33.41 0.24
Method FID $\downarrow$ CLIP $\uparrow$
No guidance 33.41 0.24
AG 26.36 0.25
CFG indicator 24.33 0.26
Model FID $\downarrow$ CLIP $\uparrow$
CLIP (0.4B) 24.33 0.26
T5-XXL (11B) 22.28 0.27
Llama3 (7B) 21.20 0.27
Train strategy FID $\downarrow$ CLIP $\uparrow$
2-stage separate training 32.55 0.24
Joint training 24.33 0.26
2-stage w/ joint finetuning 23.79 0.26
Table 3: Ablation study on Text Variational Encoder, training objective, CFG, language models, and training strategy. We conduct ablation study on our smallest model (70M), reporting zero-shot FID-10K and CLIP scores. Final settings used for CrossFlow are underlined. AG: Autoguidance. *: results without applying CFG.
Variational Encoder vs. standard encoder. Compared to a standard encoder or even adding Guassian noise like CFM [23], a Variational Encoder significantly improves the generation quality, with significant gains in the FID. This shows that forming a regularized distribution for the source domain is a crucial step for cross-modal flow matching.
Joint training vs. two-stage training. We consider three training strategies: (1) jointly training the VE and flow matching from scratch, (2) training the VE first and then training flow matching with a fixed VE, and (3) training the VE first and then training the flow matching while jointly fine-tuning VE. We observe that it is important to update the VE when training the flow matching, either through joint training from scratch, or finetuning the VE jointly with flow matching. Initializing with a pre-trained VE and then jointly training improves convergence speed by about $35\%$, but we opt to jointly train both models from scratch on account of the simplicity, and for fair comparisons with baselines.
CFG indicator. We evaluate the performance of our model when leveraging our proposed CFG indicator techinuqe. We also evaluate Autoguidance (AG) [38], which utilizes two models for inference – we use an under-trained version of the same model as the bad model, while using a grid-search to find the best under-trained checkpoint. While AG improves FID and also image-text CLIP alignment slightly, our CFG indicator works better than AG in terms of both FID and CLIP alignment while only using a single model trained with standard CFG settings. Qualitatively, our approach produces much higher fidelity images compared to both alternatives, as shared in Tab. 9.
Text VE loss. We explore reconstruction and contrastive objectives for the encoder loss $L_{Enc}$ when training the text VE. We find that a contrastive loss which promotes semantic understanding yields significantly better performance than reconstruction loss on the input text embeddings. Furthermore, image-text contrastive loss slightly outperforms text-text contrastive loss.
Effect of different language models. We evaluate CrossFlow with various language models trained with different objectives. Specifically, we evaluate CLIP [66] (contrastive image-text), T5-XXL’s encoder [67] (encoder-decoder), Llama3-7B [19] (decoder-only). We use $77$ tokens for all language models, resulting in text embeddings of size $77\times 768$, $77\times 4096$, $77\times 4096$, respectively. We train a separate Text VE for each language model, projecting the text embeddings into the target image latent shape ($4\times 32\times 32$). CrossFlow works well with all language models regardless of their training objectives and embedding sizes. As expected, our performance improves with better text representations. Due to compute restrictions however, we use the light-weight CLIP model for our main experiments.
5.3 CrossFlow for Various Tasks
We further evaluate CrossFlow on three distinct tasks that involve cross-modal / intra-modal evolution. We present the main results and key findings here, while additional details and qualitative results can be found in the Appendix.
Image to text (captioning). We first consider the task of image captioning. To achieve this, we train a new Text Variational Encoder on the captioning dataset to extract text latents from text tokens, and a separate text decoder with a reconstruction loss to convert text latents back into tokens. CrossFlow is then trained to map from the image latent space to the text latent space. Following previous work, we use the Karpathy split [37] of COCO dataset [49] for training and testing, and report results in Tab. 4. CrossFlow enables direct evolution from image space to text space for image captioning, achieving state-of-the-art performance.
Method B@4 $\uparrow$ M $\uparrow$ R $\uparrow$ C $\uparrow$ S $\uparrow$
MNIC [24] 30.9 27.5 55.6 108.1 21.0
MIR [43] 32.5 27.2 - 109.5 20.6
NAIC-CMAL [28] 35.3 27.3 56.9 115.5 20.8
SATIC [96] 32.9 27.0 - 111.0 20.5
SCD-Net [58] 37.3 28.1 58.0 118.0 21.6
CrossFlow (Ours) 36.4 27.8 57.1 116.2 20.4
Table 4: Image captioning on COCO Karpathy split. CrossFlow directly evolves from image to text, achieving comparable performance to state-of-the-art models on image captioning. For a fair comparison, we only consider non-autoregressive methods that are trained without CIDEr optimization.
Image to depth (depth estimation). For monocular depth estimation, we train CrossFlow in pixel space. Specifically, we use a recontruction loss to train the Image Variational Encoder to map the original image into the shape of a depth map, followed by the flow matching model which generates the final depth maps. We train and evaluate our model on KITTI [25] (Eigen split [21]) and NYUv2 [78] (official split) for outdoor and indoor scenarios, respectively. As shown in Tab. 5, our model achieves comparable performance to state-of-the-art methods on both datasets. Notably, DiffusionDepth [18] utilizes Swin Transformer [56] and specific designs such as Multi-Scale Aggregation and Monocular Conditioned Denoising Block. In contrast, our model achieves similar performance without any additional enhancements, demonstrating the efficiency and effectiveness of CrossFlow in mapping from images to depth.
Method KITTI NYUv2
AbsRel ($\downarrow$) $\boldsymbol{\delta_{1}}$ ($\uparrow$) AbsRel ($\downarrow$) $\boldsymbol{\delta_{1}}$ ($\uparrow$)
TransDepth [89] 0.064 0.956 0.106 0.900
AdaBins [6] 0.058 0.964 0.103 0.903
DepthFormer [45] 0.052 0.975 0.096 0.921
BinsFormer [46] 0.052 0.974 0.094 0.925
DiffusionDepth [18] 0.050 0.977 0.085 0.939
CrossFlow (Ours) 0.053 0.973 0.094 0.928
Table 5: Monocular depth estimation on KITTI and NYUv2. CrossFlow enables direct mapping from image to depth, achieving comparable performance to state-of-the-art models.
Low-resolution to high-resolution (super-resolution). We compare CrossFlow with the standard super-resolution method using flow matching, which involves upsampling the low-resolution image and then concatenating it with input noise as conditioning before feeding it into the neural network. In contrast, our method directly evolves the upsampled low-resolution image into a high-resolution image, without additional concatenation conditioning. We also compare against SR3 [75] which uses diffusion models for super-resolution. Following previous work [75, 50], we train and evaluate our model on ImageNet [16] for the task of $64\times 64\rightarrow 256\times 256$ super-resolution, and provide results in Tab. 6. Our method achieves better results compared to the standard flow matching baseline and SR3, indicating that CrossFlow can also effectively evolve between similar distributions while achieving superior performance.
Method FID $\downarrow$ IS $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
Reference 1.9 240.8 - -
Regression 15.2 121.1 27.9 0.801
SR3 [75] 5.2 180.1 26.4 0.762
Flow Matching [50] 3.4 200.8 24.7 0.747
CrossFlow (Ours) 3.0 207.2 25.6 0.764
Table 6: Image super-resolution on the ImageNet validation set. Compared with standard SR method with flow matching, our direct mapping method achieves better performance.
6 Conclusion
In this paper, we proposed CrossFlow, a simple and general framework for cross-modal flow matching that works well across a variety of tasks without requiring task specific architectural modifications. It outperforms conventional flow matching, while also enabling new capabilities such as latent arithmetic. We showcase that CrossFlow is a promising approach for the future thanks to its better scalablity. We hope our approach helps pave the way towards further research and applications of cross-modal flow matching.
Acknowledgements. We sincerely appreciate Ricky Chen and Saketh Rambhatla for their valuable discussions.
Appendix
In the appendix, we provide additional information as listed below:
  • Sec. A. Method details
    • Sec. A.1. Loss function for text-to-image generation
    • Sec. A.2. Experimental details for various tasks
  • Sec. B. Additional experimental results
    • Sec. B.1. GenEval performance for text-to-image
    • Sec. B.2. Zero-shot depth estimation
    • Sec. B.3. Image super-resolution
    • Sec. 9. Ablations on Text VE and CFG indicator
  • Sec. C. Additional qualitative examples
    • Fig. 9. Text-to-image generation
    • Fig. 10,  11. Interpolation in latent space
    • Fig. 12. Arithmetic in latent space
Appendix A Method Details
A.1 Loss Function for T2I Generation
We jointly train the Text Variational Encoder with the flow matching model using the following training objective:
$L$ $=L_{FM}+L_{Enc}+\lambda L_{KL}$
$=\texttt{MSE}(v_{\theta}(z_{t},t),\hat{v})+\texttt{CLIP}(z_{0},\hat{z})$
$\qquad+\lambda\texttt{KL}(\mathcal{N}(\bar{\mu}_{z_{0}},\bar{\sigma}_{z_{0}}^{2})||\mathcal{N}(0,1))$ (5)
where $\lambda$ is the weight of KL-divergence loss. For the flow matching loss $L_{FM}$, we follow previous work [50] and compute the MSE loss between the predicted velocity $v_{\theta}(z_{t},t)$ at time-step $t$ and the ground-truth velocity $\hat{v}$. To train the Text Variational Encoder, we adopt a CLIP contrastive loss. Specifically, given a batch of $N$ text and image pairs, we use our Text Variational Encoder to obtain text latents $z_{0}$, and an image encoder to extract image features $\hat{z}$. Then, we compute the cosine similarity between all pairs of $z_{0}$ and $\hat{z}$ in the batch, resulting in a similarity matrix $S$, where each element $s_{ij}$ represents the cosine similarity between the $i^{th}$ $z_{0}$ and $j^{th}$ $\hat{z}$. The similarity scores are then scaled by a temperature parameter $\tau$ (a learnable parameter), denoted as $\text{logit}s_{ij}=s_{ij}/\tau$. After that, a symmetric cross-entropy loss over the similarity scores is computed:
$L_{\text{I2T}}=-\frac{1}{N}\sum_{i=1}^{N}\text{log}\frac{\text{exp}(\text{logit}s_{ii})}{\sum_{j=1}^{N}\text{exp}(\text{logit}s_{ij})}$ (6)
$L_{\text{T2I}}=-\frac{1}{N}\sum_{i=1}^{N}\text{log}\frac{\text{exp}(\text{logit}s_{ii})}{\sum_{j=1}^{N}\text{exp}(\text{logit}s_{ji})}$ (7)
Finally, we compute the average of these two components to obtain the CLIP loss, which is then used to update our Text Variational Encoder:
$L_{Enc}=\texttt{CLIP}(z_{0},\hat{z})=\frac{1}{2}(L_{\text{I2T}}+L_{\text{T2I}})$ (8)
For the KL loss $L_{KL}$, we adopt the original KL divergence loss [42] with $\lambda=1\times 10^{-4}$.
A.2 Experimental Details for Various Tasks
Image captioning. We conduct our experiments on the popular Karpathy split [37] of COCO dataset [49], which contains $113,287$ images for training, $5,000$ images for validation, and $5,000$ image for testing. We train our model with $351$M parameters on the training split for 100 epochs, using a batch size of $256$ and a base learning rate of $2\times 10^{-4}$ with $5$ warm-up epochs. Following the standard evaluation setup, we compare the performance over five metrics: BLEU@4 [61] (B@4), METEOR [4] (M), ROUGE [48] (R), CIDEr [87] (C), and SPICE [3] (S).
Monocular depth estimation. We consider KITTI [25] and NYUv2 [78] for outdoor and indoor depth estimation. For KITTI, we use the Eigen split [21], consisting of $23,488$ training images and $697$ testing images. For NYUv2, we adopt the official split, which contains $24,231$ training images and $654$ testing images. We train our model with $527$M parameters on the corresponding training splits for $50$ epochs. We use a batch size of $64$, and decay the learning rate from $1\times 10^{-4}$ to $1\times 10^{-8}$ with cosine annealing.
Image super-resolution. We consider natural image super-resolution, training our model on ImageNet 1K [16] for the task of $64\times 64\rightarrow 256\times 256$ super-resolution. We use the dev split for evaluation. During training, we preprocess the images by removing those where the shorter side is less than $256$ pixels. The remaining images are then centrally cropped and resized to $256\times 256$. The low-resolution images are then generated by downsampling the $256\times 256$ images using bicubic interpolation with anti-aliasing enabled. For a fair comparison with SR3 [75], we train our CrossFlow with $505$M parameters (compared to $625$M parameters in SR3). Our model is trained for $1$M training steps with a batch size of $512$ and a learning rate of $1\times 10^{-4}$, including $5,000$ warm-up steps.
Appendix B Additional Experimental Results
B.1 GenEval Performance
Method Overall Single Two Counting Colors Position Attribute
Object Object binding
DALL·E 2 [69] 0.52 0.94 0.66 0.49 0.77 0.10 0.19
LDMv1.5 [73] 0.43 0.97 0.38 0.35 0.76 0.04 0.06
LDMv2.1 [73] 0.50 0.98 0.51 0.44 0.85 0.07 0.17
LDM-XL [63] 0.55 0.98 0.74 0.39 0.85 0.15 0.23
PixArt-$\alpha$ [10] 0.48 0.98 0.50 0.44 0.80 0.08 0.07
LDMv3 ($512^{2}$[22] 0.68 0.98 0.84 0.66 0.74 0.40 0.43
CrossFlow 0.55 0.98 0.72 0.39 0.82 0.18 0.21
Table 7: GenEval comparisons. Our model achieves comparable performance to state-of-the-art models such as LDM-XL and DALL·E 2, suggesting that CrossFlow is a simple and promising direction for state-of-the-art media generation.
To compare with recent text-to-image models on GenEval, we report the overall score and task-specific scores in Tab. 7. Our model achieves comparable performance to state-of-the-art models such as LDMv2.1 [73], LDM-XL [63], and DALL·E 2 [69]. This demonstrates that directly evolving from text space to image space with our approach is a simple and effective solution for text-to-image generation, indicating a novel and promising direction for state-of-the-art media generation.
B.2 Zero-shot Depth Estimation
Method # Training samples KITTI NYUv2 ETH3D ScanNet DIODE
AbsRel $\downarrow$ $\boldsymbol{\delta_{1}}$uparrow AbsRel $\downarrow$ $\boldsymbol{\delta_{1}}$uparrow AbsRel $\downarrow$ $\boldsymbol{\delta_{1}}$uparrow AbsRel $\downarrow$ $\boldsymbol{\delta_{1}}$uparrow AbsRel $\downarrow$ $\boldsymbol{\delta_{1}}$uparrow
DiverseDepth [91] 320K 0.117 0.875 0.190 0.704 0.228 0.694 0.109 0.882 0.376 0.631
MiDaS [70] 2M 0.111 0.885 0.236 0.630 0.184 0.752 0.121 0.846 0.332 0.715
LeReS [92] 300K + 54K 0.090 0.916 0.149 0.784 0.171 0.777 0.091 0.917 0.271 0.766
Omnidata [20] 11.9M + 310K 0.074 0.945 0.149 0.835 0.166 0.778 0.075 0.936 0.339 0.742
HDN [94] 300K 0.069 0.948 0.115 0.867 0.121 0.833 0.080 0.939 0.246 0.780
DPT [71] 1.2M + 188K 0.098 0.903 0.100 0.901 0.078 0.946 0.082 0.934 0.182 0.758
Marigold [39] 74K 0.060 0.959 0.105 0.904 0.071 0.951 0.069 0.945 0.310 0.772
CrossFlow (Ours) 74K 0.062 0.956 0.103 0.908 0.085 0.944 0.068 0.942 0.270 0.768
Table 8: Zero-shot depth estimation. Baseline results are reported by Marigold [39]. We follow Marigold and train our CrossFlow on the same datasets, i.e., Hypersim [72] and Virtual KITTI [9]. We highlight the best, second best, and third best entries. With just a unified framework, CrossFlow achieves comparable or even superior performance on complex zero-shot depth estimation, demonstrating the general-purpose nature of CrossFlow on various cross-modal tasks.
Figure 6: Qualitative examples for zero-shot depth estimation. The input images in the first two rows are from the NYUv2 dataset, while the input images in the last row were generated by our T2I model. Our model provides robust zero-shot depth estimation across domains, whether indoor or outdoor, synthetic or real.
We also evaluate CrossFlow on zero-shot depth estimation. Following Marigold [39], we train our model on Hypersim [72] and Virtual KITTI [9], and evaluate our model on 5 real datasets that are not seen during training: KITTI [25], NYUv2 [78], ETH3D [76], ScanNet [13], and DIODE [85]. We follow Marigold [39] to prepare the training and testing data. Our model with $527$M parameters is trained for 150K training steps, with a batch size of 512 and a learning rate of $1\times 10^{-4}$ with $5,000$ warm-up steps. The results are reported in Tab. 8. Qualitative examples are provided in Fig. 6. Without specific design, CrossFlow achieves comparable or even superior performance compared to state-of-the-art methods, demonstrating the general-purpose nature of our approach on various cross-modal tasks.
B.3 Image Super-resolution
We provide qualitative examples for image super-resolution in Fig. 7. Unlike traditional methods, which typically evolve from Gaussian noise and rely on concatenating upsampled low-resolution images as conditioning, our approach takes a more direct route: we demonstrate that it is possible to evolve a low-resolution image directly into a high-resolution image, eliminating the need for additional concatenation conditioning.
Figure 7: Qualitative examples for image super-resolution.
B.4 Ablation Study
Text encoder Recon. accuracy (%)
Text Encoder ($1\times 1024$) 95.12
Text Variational Encoder ($1\times 1024$) 94.53
Table 9: Ablation on text compression. Both text encoder and Text Variational Encoder preserve most of the input information, despite the large compression ratio ($77\times 768\rightarrow 1\times 1024$, 14.4×14.4$\times$14.4 ×).
Text compression. In this section, we show that we can compress the input text embedding $x\in\mathbb{R}^{n\times d}$ into $z_{0}\in\mathbb{R}^{h\times w\times c}$ (e.g., $77\times 768$ CLIP tokens to $4\times 32\times 32$ latents for $256$px generation, 14.4×14.4$\times$14.4 × compression) with a standard encoder or the proposed Variational Encoder while preserve most of the input information. We report the per-token reconstruction accuracy, computed by cosine similarity, in Tab. 9. The results show that both methods are effective at preserving the input information, achieving high reconstruction accuracy despite a large compression ratio.
Figure 8: Ablation on CFG with indicator. The first two columns show the images generated when the indicator $01_{c}=$ 0 (for unconditional generation) and $11_{c}=$ 1 (for conditional generation), demonstrating that CrossFlow can still perform unconditional generation with the help of the indicator, thereby allowing for the use of standard CFG. We then demonstrate the improvement provided by CFG (middle five columns) and compare it with Autoguidance (last two columns). Prompts used to generate the images: ‘a corgi wearing a red hat in the park’,‘a cat playing chess’,‘a pair of headphones on a guitar’,‘a horse in a red car’
CFG indicator. In Fig. 8, we study the effect of our CFG with indicator, and then compare our approach with Autoguidance [38]. The left two columns show the images generated when the indicator $01_{c}=$ 0 (for unconditional generation) and $11_{c}=$ 1 (for conditional generation). It shows that despite generating an image by directly evolving from the text space into the image space without explicit conditioning, our model can still perform unconditional generation with the help of the indicator. This allows our model to support standard CFG. Then, in the middle five columns, we show the images generated with different CFG scaling factors. Similar to the standard flow matching model, the CFG can significantly improve the image quality. Finally, in the last two columns, we compare our CFG with indicator to Autoguidance, using the same scaling factor. Like our approach, Autoguidance also enables low-temperature sampling for models without explicit conditioning. We observe that our CFG with indicator produces higher-fidelity images compared to Autoguidance.
Appendix C Additional Qualitative Examples
We provide additional qualitative examples for text-to-image generation here. Specifically, we first provide $512\times 512$ images generated by our CrossFlow in Fig. 9. Next, we provide more examples for linear interpolation in latent space (Fig. 10 and Fig. 11) and arithmetic operation in latent space (Fig. 12).
Figure 9: Qualitative examples for text-to-image with CrossFlow.
Figure 10: Linear interpolation in latent space. We show images generated by linear interpolation between two text latents (i.e., interpolation between $z_{0}$). Images generated by the first and second text latents are provided in the top-left and bottom-right corners.
Figure 11: Linear interpolation in latent space. We show images generated by linear interpolation between two text latents (i.e., interpolation between $z_{0}$). Images generated by the first and second text latents are provided in the top-left and bottom-right corners.
Figure 12: Arithmetic in text latent space. We map the text into the text latent space, perform arithmetic operations to obtain new latent representation, and use the resulting representation to generate the image. Latent $z_{0}$ used to generate each image is provided at the bottom.